The Largo Publishing House uses 400 printers and 200 printing presses to produce books. A printer’s wage rate is $20, and the price of a printing press is $5,000. The last printer added 20 books to total output, while the last press added 1,000 books to total output. Is the publishing house making the optimal input choice? Why or why not? If not, how should the manager of Largo Publishing House adjust input usage?
Answer #1

Marginal product of printer = 20

Marginal product of the last press = 1000

MP per unit input price (printer) = 20/20 = 1

MP per unit input price (printing press) = 1000/5000 = 0.2

The publishing house is not making the optimal input choice because MP per unit input price is not the same for the given inputs. It must employ more printers and less printing presses to produce books.